OpenStudy (anonymous):

Find the formula for the 17th term of the sequence in which a1 = -3 and the common ratio is -2

OpenStudy (anonymous):

An=a1 r^n-1 A17th =(-2)^17-1 =(-2)^16 ANSWER...YES OR NO? =65,536 ?